OpenStudy (anonymous):

the area of a rectangular barnyard is given by the trinomial 6x^2+7x-20

OpenStudy (anonymous):

god girl

OpenStudy (anonymous):

@GODGIRL

OpenStudy (anonymous):

Whats ur answer choices?

OpenStudy (anonymous):

@crazyinlove

OpenStudy (anonymous):

2x-5 and 3x-4 -2x-5 and -3x-4 2x-5 and 3x-4 2x-5 and 3x-4

OpenStudy (anonymous):

kk jus a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

uhmm sorry but i have no idea on how to figure this out maybe @AccessDenied or @mathmale could help . :) Sorry again! wish i could help :/

OpenStudy (anonymous):

thank u for trying

OpenStudy (accessdenied):

Area for rectangle is base times height. So if we can find two factors of 6x^2+7x-20, those would be our base and height. Did you have any ideas on how to begin factoring it?

OpenStudy (anonymous):

no not really

OpenStudy (accessdenied):

I believe from ax^2 + bx + c, we start by looking for factors of a*c that add up to b. So a*c = 6*-20. One factor will be negative, and we have to add the two factors to get: b = 7 -120 = 15*-8 <-- these two factors seem to add up to 7, so we rewrite: 6x^2 + 7x - 20 = 6x^2 + (15x - 8x) - 20 = (6x^2 + 15x) + (-8x - 20) So that we can factor by grouping. Factor each individual group until we get a similar factor in each, then pull that factor out.

OpenStudy (accessdenied):

The first group we could pull out a 3x. 6x^2 + 15x = 3x (2x + 5). The second group, we could take out -4: -8x - 20 = -4 (2x + 5) So that both factors hold a (2x + 5), we pull it out: (2x + 5)(3x - 4)

OpenStudy (anonymous):

thats actually what i got but that wasn't a answer choice so i was like........stuck basically

OpenStudy (accessdenied):

I'm guessing one of the last answer choices is written incorrectly. (They're both the same, one must have a + sign)

OpenStudy (accessdenied):

Actually, now that I look at it again... the first, third AND fourth are the same. o.o

OpenStudy (anonymous):

Yeah so @crazyinlove do any of ur answer choices have (2x+5)(3x-4)

OpenStudy (anonymous):

ok thank u

OpenStudy (accessdenied):

Alternatively, you could work in reverse: just expand out each answer choice and see which one gives you the correct area. That may or may not take more time if you don't remember the factoring method.